### v1

@@ 3.1 Cluster-tree block partitioning
         BuildPartition$(\boldsymbol{\tau}^{\prime},\boldsymbol{\sigma}^{\prime})$
@@ 3.2 Adaptive cross approximation (ACA)
where ||$\cdot$|| ⋅ | denotes the determinant of a matrix. Notably, unlike standard $\mathcal{H}$-matrix definitions where elements are scalars, each element in our matrix is a $3\times 3$ submatrix, as previously explained. In ACA, the iteration process starts with a randomly selected initial column index, $j_{1}$, and terminates when the magnitude of a new update is sufficiently small, i.e.,
@@
where $\|\cdot\|_{F}$ denotes the Frobenius norm; $\zeta$ is the preset tolerance. The complete procedure of ACA employed in this work is presented in Algorithm 2. Some notations used therein are explained below. Rows and columns of a matrix $\mathbf{A}$ are denoted as $\mathbf{A}(I,:)$ and $\mathbf{A}(:,I)$, respectively, where $I$ is the index set. Vertical and horizontal concatenations of $\mathbf{A}$ and $\mathbf{B}$ are represented by $[\mathbf{A};\mathbf{B}]$ and $[\mathbf{A},\mathbf{B}]$, respectively. As the $i$-th element of $\mathbf{c}$ is denoted by $\mathbf{c}(i)$ (a $3\times 3$ submatrix), the inverse of $\mathbf{c}(i)$ is written as $[\mathbf{c}(i)]^{-1}$.
@@
[…]
@@ 4.2 Assessment of complexity and scalability
Figure 8: Number of query vs. $N$ under different $\zeta$, with $C_{\text{leaf}}=50$. The black dashed line denotes the scaling of $O(N\log N)$.
@@
Figure 9: Number of query vs. $N$ under different $C_{\text{leaf}}$, with $\zeta=0.1$. The black dashed line denotes the scaling of $O(N\log N)$.
@@
(a) Wall time vs $N$ on different numbers of GPUs. The black dashed line denotes the scaling of $O(N\log N)$.
@@
Figure 10: Scalability test of wall time with respect to $N$ (number of particles), ranging from $10^{5}$ to $10^{7}$, and $N_{\text{GPU}}$ (number of GPUs), ranging from 1 to 8 on a single node. Here, $\zeta=0.1$ and $C_{\text{leaf}}=50$
@@ 4.3.1 Suspension drops
In the first case, we simulated the interaction of two large suspension drops while sedimenting subject to gravity. Composed of a cloud of particles, each suspension drop initially formed a sphere with a radius of 175, with the volume fraction of particles as 10.0%. We constructed the initial spherical suspension drop by first generating a primitive cubic lattice structure, then discarding the particles lying outside the desired spherical domain, and finally slightly perturbing the position of each particle randomly to emulate a real suspension drop more accurately. Each suspension drop is composed of 523,305 particles, and hence the entire system contains totally 1,046,610 (more than 1 million) particles. The gravitational force acting on each particle is given by $\mathbf{F}=[0,0,-1]$, directed along $-z$ direction. For this simulation, the time step was set to $\Delta t=0.01$, and the $\mathcal{H}$-matrix parameters were set to $\zeta=0.1$ and $C_{\text{leaf}}=50$.
@@ 4.3.2 Flexible filaments
where $\mathbf{r}_{i\to i+1}=\mathbf{X}_{i+1}-\mathbf{X}_{i}$; $l_{0}$ is the rest length and equal to $r_{b}$, the initial separation distance set between any two adjacent beads; and $k_{t}$ is the tension stiffness coefficient and was set to $k_{t}=20$ in this work. This bonding force is added to the associated particles $i$ and $i+1$ by: $\mathbf{F}_{i}\leftarrow\mathbf{F}_{i}+\mathbf{f}^{\text{tension}}_{i\to i+1}$ and $\mathbf{F}_{i+1}\leftarrow\mathbf{F}_{i+1}-\mathbf{f}^{\text{tension}}_{i\to i+1}$, respectively. The bending force acts on three consecutive beads $(i-1,i,i+1)$ along each filament chain and is derived from the COS bending potential [49, 52]:

### v2

@@ 3.1 Cluster-tree block partitioning
[…]
@@ 3.2 Adaptive cross approximation (ACA)
where $|\cdot|$ denotes the determinant of a matrix. Notably, unlike standard $\mathcal{H}$-matrix definitions where elements are scalars, each element in our matrix is a $3\times 3$ submatrix, as previously explained. In ACA, the iteration process starts with a randomly selected initial column index, $j_{1}$, and terminates when the magnitude of a new update is sufficiently small, i.e.,
@@
where $\|\cdot\|_{F}$ denotes the Frobenius norm; $\zeta$ is the preset tolerance. The complete procedure of ACA employed in this work is presented in Algorithm 2. Some notations used therein are explained below. Rows and columns of a matrix $\mathbf{A}$ are denoted as $\mathbf{A}(I,:)$ and $\mathbf{A}(:,I)$, respectively, where $I$ is the index set. Vertical and horizontal concatenations of $\mathbf{A}$ and $\mathbf{B}$ are represented by $\mathbf{A};\mathbf{B}]$ and $\mathbf{A},\mathbf{B}]$, respectively. As the $i$-th element of $\mathbf{c}$ is denoted by $\mathbf{c}(i)$ (a $3\times 3$ submatrix), the inverse of $\mathbf{c}(i)$ is written as $\mathbf{c}(i)]^{-1}$.
@@
13:  $\mathbf{C}=[\mathbf{C},\mathbf{c}_{k}]$, $\mathbf{Q}=[\mathbf{Q};\mathbf{q}^{T}_{k}]$.
@@ 4.2 Assessment of complexity and scalability
Figure 8: Number of query vs. NN under different $\zeta$, with $C_{\text{leaf}}=50$. The black dashed line denotes the scaling of $O(N\log N)$.
@@
Figure 9: Number of query vs. NN under different $C_{\text{leaf}}$, with $\zeta=0.1$. The black dashed line denotes the scaling of $O(N\log N)$.
@@
(a) Wall time vs NN on different numbers of GPUs. The black dashed line denotes the scaling of $O(N\log N)$.
@@
Figure 10: Scalability test of wall time with respect to NN (number of particles), ranging from $10^{5}$ to $10^{7}$, and $N_{\text{GPU}}$ (number of GPUs), ranging from 1 to 8 on a single node. Here, $\zeta=0.1$ and $C_{\text{leaf}}=50$
@@ 4.3.1 Suspension drops
In the first case, we simulated the interaction of two large suspension drops while sedimenting subject to gravity. Composed of a cloud of particles, each suspension drop initially formed a sphere with a radius of 175, with the volume fraction of particles as 10.0%. We constructed the initial spherical suspension drop by first generating a primitive cubic lattice structure, then discarding the particles lying outside the desired spherical domain, and finally slightly perturbing the position of each particle randomly to emulate a real suspension drop more accurately. Each suspension drop is composed of 523,305 particles, and hence the entire system contains totally 1,046,610 (more than 1 million) particles. The gravitational force acting on each particle is given by $\mathbf{F}=[0,0,-1]$, directed along z-z direction. For this simulation, the time step was set to $\Delta t=0.01$, and the $\mathcal{H}$-matrix parameters were set to $\zeta=0.1$ and $C_{\text{leaf}}=50$.
@@ 4.3.2 Flexible filaments
where $\mathbf{r}_{i\to i+1}=\mathbf{X}_{i+1}-\mathbf{X}_{i}$; $l_{0}$ is the rest length and equal to $r_{b}$, the initial separation distance set between any two adjacent beads; and $k_{t}$ is the tension stiffness coefficient and was set to $k_{t}=20$ in this work. This bonding force is added to the associated particles $i$ and $i+1$ by: $\mathbf{F}_{i}\leftarrow\mathbf{F}_{i}+\mathbf{f}^{\text{tension}}_{i\to i+1}$ and $\mathbf{F}_{i+1}\leftarrow\mathbf{F}_{i+1}-\mathbf{f}^{\text{tension}}_{i\to i+1}$, respectively. The bending force acts on three consecutive beads (i1,i,i+1)(i-1,i,i+1) along each filament chain and is derived from the COS bending potential [57, 60]:
@@ Acknowledgments
Data Availability
The codes and examples that support the findings of this study are available in the link: https://github.com/Pan-Group-UW-Madison/hignn.